OpenStudy (anonymous):

HELPP solve the quadratic equation for x by factoring. x^2 - 13x = 0

OpenStudy (anonymous):

\[x^{2} - 13x = 0 \]

OpenStudy (wolf1728):

x² = 13x x² / x = 13 x = 13

OpenStudy (anonymous):

Wolf..you can't really do that... x^2 = 13x Subtract 13x from both sides, you get x^2 - 13x = 0 Now factor, you get x(x - 13) = 0 Set each factor equal to 0, you get x = 0 and x = 13. You cannot divide both sides by x, or you lose a solution of x = 0.

OpenStudy (wolf1728):

Well I guess I lost a solution. Actually I wasn't dividing both sides by x. I was reducing the fraction x²/x multiply top and bottom by 1/x results in x.

OpenStudy (anonymous):

thank you guys :)